A ball of radius 0.200 m rolls with a constant linear speed of 3.01 m/s along a horizontal table. The ball rolls off the edge and falls a vertical distance of 2.20 m before hitting the floor. What is the angular displacement of the ball while the ball is in the air?
Answer #1

Solution,

Given,

Angular velocity, w = v/r = 3.01/0.20 = 15.05 rad/s

Using kinematic equation

Y = 0.5 gt^2

Time, t = sqrt(2y/g) = sqrt(2 x 2.2/9.8) = 0.67 sec

Angular displacement, theta = wt = 15.05 x 0.67 = 10.08 rad

Comment in case any doubt please rate my answer ...